# SLOPE and Designing Robust Studies for Generalization

### v1

@@ 2.1 Goal and Key Assumptions in Generalization
$Q_{X}$ is absolute continuous with respect to $P_{X}$.
@@
$Q_{O\mid X}(\cdot\mid X)$ is absolute continuous with respect to $P_{O\mid X}(\cdot\mid X)$ and the Radon-Nikodym derivative satisfies $dQ_{O\mid X}(O,X)/dP_{O\mid X}(O,X)=1$ almost everywhere $P_{O\mid X}\times Q_{X}$.
@@
We note that Assumption (2) is equivalent to equation (1) stated in Section 1. Further, when the density of $P_{O\mid X}$ exists and the density of $Q_{O\mid X}$ is assumed to exist, then Assumption 2 can be re-formulated as a condition on the density ratio, i.e., $f_{Q_{O\mid X}}(O,X)/f_{P_{O\mid X}}(O,X)=1$ almost everywhere $P_{O\mid X}\times Q_{X}$. For presentation purposes, we keep this density ratio representation in later sections, while keeping in mind that this can be formulated as a Radon-Nikodym derivative as in Assumption 2.
Assumption 1 states that the support of $Q_{X}$ is within the support of $P_{X}$. Assumption 2 enables the identification of $\psi(Q_{O,X})$ by replacing the unidentified $Q_{O\mid X}$ with the identified $P_{O\mid X}$. Under Assumptions 1 and 2, the target estimand can be identified as $\psi(Q_{X}\times P_{O\mid X})$. We provide some examples of target estimands and their identification below.
@@
Suppose we are interested in the mean of $O$ in the target distribution, denoted as $\psi^{{\textnormal{mean}}}(Q_{O})=\textnormal{E}_{Q_{O}}(O)$. Under Assumptions 1 and 2, the mean is identified via
$\psi^{{\textnormal{mean}}}(Q_{O})=\textnormal{E}_{Q_{O}}(O)=$ $\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{Q_{O\mid X}}(O)\right]=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O)\right].$
@@
Suppose we are interested in the median of $O$ in the target distribution, denoted as $\psi^{{\textnormal{med}}}(Q_{O})=F_{Q_{O}}^{-1}(1/2)$, and $O$ is continuous. Under Assumptions 1 and 2, the median is identified as the solution to the following equation:
@@
Example 3 (Z-Functionals).
Suppose the target estimand $\psi(Q_{O,X})$ is defined as the solution to the equation
@@
where $s(O,X,\cdot)$ is a user-specified function, usually a score function, of the same dimension with $\psi$. Under Assumptions 1 and 2, $\psi$ is identified as the solution to the following equation:
$0=\textnormal{E}_{Q_{O,X}}\{s(O,X,\psi)\}=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{Q_{O\mid X}}\left\{s(O,X,\psi(Q_{O\mid X}\times Q_{X}))\right\}\right]=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\left\{s(O,X,\psi(Q_{O,X}^{0}))\right\}\right].$
@@
Suppose a randomized experiment has been conducted in the source population to measure the causal effect of treatment on an outcome. Let $Y(a)$ be the potential outcome if a study unit was assigned to treatment level $a\in\{0,1\}$ and $X$ be the pre-treatment covariates. The goal is to use this randomized experiment to learn about the causal effect in a new, target population based on the target population’s covariates $X$. Under Assumptions 1 and 2, the mean and the median of the potential outcome $Y(a)$ are identified via
$\psi^{{\textnormal{mean}}}(Q_{Y(a)})=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{Y(A)\mid X}}\{Y(A)\}\right],\text{ and }\quad{}\frac{1}{2}=\int\int_{-\infty}^{\psi^{{\textnormal{med}}}}dP_{Y(A)\mid X}dQ_{X}.$
We note that both equations involve potential outcomes (i.e., $P_{Y(a)\mid X}$), which is identified under a randomized experiment; see Section D in the Supplement for details.
@@ 2.2 Model for Sensitivity Analysis of Conditional Exchangeability
The identification of the target estimand depends on Assumptions 1 (i.e., overlap) and 2 (i.e., conditional exchangeability). While Assumption 1 can be empirically verified with observed data (i.e., $X$ from $P_{X}$ and $Q_{X}$), Assumption 2 cannot be verified since it concerns unobserved data, specifically $O$ in the target population $Q$. Faced with the inability to check the assumption, a sensitivity analysis seeks to study how conclusions from a study may change if Assumption 2 is violated by a certain amount. In this section, we describe a model for sensitivity analysis, which motivates how we define SLOPE.
Suppose conditional exchangeability (i.e., Assumption 2) no longer holds and instead, the distribution $Q_{O\mid X}$ is shifted away from $P_{O\mid X}$ by an exponential tilting term. Specifically, we consider Robins’ sensitivity model (Robins et al.,, 2000) where $Q_{O\mid X}^{\gamma}$ is the exponentially tilted distribution of $Q_{O\mid X}$. Suppose $Q_{O\mid X}^{\gamma}(\cdot\mid X)$ is absolute continuous with respect to $P_{O\mid X}(\cdot\mid X)$ almost everywhere $Q_{X}$ and their corresponding densities $f_{Q_{O\mid X}^{\gamma}}(O,X)$ and $f_{P_{O\mid X}}(O,X)$ satisfy
@@
$\dfrac{f_{Q_{O\mid X}^{\gamma}}(O,X)}{f_{P_{O\mid X}}(O,X)}\propto\exp(\gamma\cdot O)\text{, almost everywhere }P_{O\mid X}\times Q_{X}.$ (3)
@@
The notation $\propto$” means “proportional to” and $\int\exp(\gamma O)dP_{O\mid X}(O\mid X)<\infty$ almost everywhere $Q_{X}$. The term $\gamma$ is a scalar sensitivity parameter that quantifies the difference between $Q_{O\mid X}$ and $P_{O\mid X}$. If $\gamma=0$, the sensitivity model (3) reduces to Assumption 2 where the two distributions are identical. As $\gamma$ moves away from zero, Assumption 2 is violated by an amount quantified by $\gamma$.
Under Assumption 1 and the sensitivity model (3), given a value of $\gamma$, the target estimand can be identified as $\psi(Q_{O,X}^{\gamma})$ where $Q_{O,X}^{\gamma}=Q_{O\mid X}^{\gamma}\times Q_{X}$ represents the joint distribution of $Q_{O\mid X}^{\gamma}$ induced from (3) and $Q_{X}$. Since sensitivity parameters (e.g., $\gamma$) are inherently unidentifiable, in a sensitivity analysis, researchers specify a range of values for the sensitivity parameter $\gamma$ based on domain knowledge (Scharfstein et al.,, 1999; Rotnitzky et al.,, 2001; Nabi et al.,, 2024), benchmarking (Huang,, 2024) or calibration (Miao et al.,, 2024) and investigate how the target estimand changes between the case with $\gamma=0$ (i.e., $\psi(Q_{O,X}^{0})$ under conditional exchangeability) and cases where $\gamma\neq 0$ (i.e., $\psi(Q_{O,X}^{\gamma})$) with a specified range of $\gamma$ that violates conditional exchangeability.
@@
We make some remarks about the sensitivity model (3). First, model (3) was proposed by Scharfstein et al., (1999) and Robins et al., (2000) as a non-parametric (just) identified model for describing selection bias in missing data. This sensitivity model was later referred to as an exponential tilting model (Rotnitzky et al.,, 2001; Birmingham et al.,, 2003) and an extrapolation-factorization model (Linero and Daniels,, 2018). The model was also used to conduct sensitivity analysis for unmeasured confounding in causal inference (Franks et al.,, 2020; Nabi et al.,, 2024) and sensitivity analysis for transportability (Dahabreh et al.,, 2022; Miao et al.,, 2024). Second, model (3) can be reformulated as a selection model or under some assumptions, the sensitivity parameter $\gamma$ can be reparameterized to pseudo-$R^{2}$ (Franks et al.,, 2020). Third, model (3) can be extended so that $\gamma$ depends on $X$ and $O$; see Remark 7 of the Supplement. Fourth, an important caveat of using model (3) is the non-collapsibility of the model with respect to $X$ as the model implies logistic selection model; see Section 7 of Scharfstein et al., (1999).
Lastly and importantly, we choose model (3) as the model (i) posits no testable implications on the data, (ii) makes our proposed measure SLOPE tractable in terms of having a unique tangent curve (see Section 4.3), and (iii) has an analytic connection to the influence function (see Section 4.1). In general, reflecting on the discussions by Robins, (2002) and Rosenbaum, (2002), we believe there is no “optimal” model for sensitivity analysis and depending on the scientific questions at hand, some models for sensitivity analysis are more useful than others. In our case, as we show below, model (3) is useful to define SLOPE and to inform robust study designs in generalizations. A further discussion on other sensitivity models is provided in Section 4.3.
@@ 3.1 Definition and Basic Properties
Under sensitivity model (3), $\psi(Q_{O,X}^{\gamma})$ can be viewed a function that maps the sensitivity parameter $\gamma$ to the target estimand. Therefore, the fluctuation of this function along $\gamma$ characterizes how much the target estimand changes when the conditional exchangeability is violated. This observation motivates us to introduce our metric SLOPE, which measures the change in the target estimand along $\gamma$ by taking a derivative with respect to $\gamma$.
@@
The sensitivity to local perturbation from exchangeability (SLOPE) of a target functional $\psi(\cdot)$ with respect to the sensitivity model in (3) is defined as
@@
[…]
When comparing the magnitudes of SLOPE, investigators should be aware that the unit of SLOPE inherits the unit of the target estimand/functional. For instance, in Examples 1 and 2 for the mean $\psi^{{\textnormal{mean}}}$ and the median $\psi^{{\textnormal{med}}}$, the SLOPE shares the unit of $O$. While this fact is rather obvious from SLOPE’s definition since it measures the change in the target estimand and the target estimand should carry scientifically meaningful units, we remind investigators to exercise the same level of caution when interpreting SLOPE’s magnitude akin as when investigators interpret the units of slopes in other areas of statistics (e.g., in regression models).
An important property about SLOPE is that it does not depend on any estimation procedure. Instead, SLOPE measures an intrinsic, robustness property about a study design when there is violation of conditional exchangeability. Specifically, SLOPE depends on two design level quantities: (a) the target functional/estimand $\psi(\cdot)$, and (b) the target distribution under conditional exchangeability ($Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$). For scientific researchers, SLOPE reflects their answers to two questions: (a) “what quantity would you like to learn?” and (b) “what type of dataset would you use to learn the target quantity?”. Changing either (a) or (b) can lead to a different value of SLOPE, reflecting a different level of robustness of the underlying study design. Consequently, SLOPE can be used to choose which estimand (e.g., the mean or the median), target distribution (e.g., $Q_{X}$), or the source distribution (e.g., $P_{O,X}$) is less sensitive for generalization, thereby informing choices of a robust study design. See Section 3.2 for an illustration and Section 5 for real data examples. Hereinafter, we may refer to the choices of target functionals (e.g., choice between $\psi^{{\textnormal{mean}}}$ and $\psi^{{\textnormal{med}}}$) as the choices of target estimands (e.g., choice between $\psi^{{\textnormal{mean}}}(Q_{O})$ and $\psi^{{\textnormal{med}}}(Q_{O})$) when the context is clear, since the term “estimand” is more familiar to practitioners.
When communicating the meaning of SLOPE, some audience may find it useful to interpret SLOPE as a measure of the “first-order change” of the estimand when conditional exchangeability is violated. Specifically, a first-order Taylor expansion of the target estimand yields
@@
The Taylor expansion in (5) suggests that for a small $\gamma$ that is near zero, the SLOPE provides an accurate measure of the change in $\psi$ when conditional exchangeability is violated. When $\gamma$ is large in magnitude, SLOPE may still provide some intuition about the bias, with the usual limitations of first-order linear approximations. We remark, however, that SLOPE cannot identify the bias of an estimator of $\psi$ from violating conditional exchangeability since, as mentioned in Section 2.2, $\gamma$ and $Q_{O,X}^{\gamma}$ are not identified.
@@
Finally, we remark that SLOPE does not always exist for every target estimand. For example, suppose $\psi$ is the sign of the mean of $O$ in the target population. Then, the limit in Definition 1 may not be well-defined when the sign changes near $\gamma=0$. One general condition for SLOPE to exist is to satisfy the conditions for the chain rule under Hadamard differentiability (Theorem 20.9 of Van der Vaart, (2000)); see Section A of the Supplement for the precise conditions. For Z-functionals defined through (2), a sufficient condition for the existence of SLOPE is guaranteed by the following smoothness and boundndess conditions that are typical to establish consistency of Z estimators.
Condition 1 (Existence of SLOPE for Z-Functionals).
@@
(i) $\textnormal{E}_{Q_{O\mid X}^{\gamma}}[s(O,X,\psi(Q_{O,X}^{0})]$ is bounded by an integrable function almost surely $Q_{X}$ for $\gamma$ in a neighborhood of zero, and $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi(Q_{O,X}^{0})\left\{O-\mu(X)\right\}\right]$ exists; (ii) $s(O,X,\cdot)$ is differentiable almost everywhere and denote the derivative as $\dot{s}(O,X,\cdot)$; and (iii) $\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi(Q_{O,X}^{0}))\}$ exists and is nonsingular.
3.2 Example: SLOPE for the Mean and the Robustness of a Homogeneous Source Population
To gain intuition, we illustrate the SLOPE when the target estimand is the mean, i.e., $\psi^{{\textnormal{mean}}}$ in Example 1 with $\psi^{\textnormal{mean}}(Q_{O,X})=\textnormal{E}_{Q_{O}}(O)$. Given $\gamma$, sensitivity model (3) implies
$\psi^{{\textnormal{mean}}}(Q_{O,X}^{\gamma})=\textnormal{E}_{Q_{X}}\left[\dfrac{\textnormal{E}_{P_{O\mid X}}\left\{O\exp(\gamma O)\mid X\right\}}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\mid X\right\}}\right].$

### v2

@@ 2.1 Goal in Generalization and Key Assumptions
$Q_{X}$ is absolutely continuous with respect to $P_{X}$.
@@
Equation (1) holds.
@@
When the densities of $P_{O\mid X}$ and $Q_{O\mid X}$ exist with respect to a common measure (e.g., the Lebesgue measure or the counting measure), Assumption 2 can be re-formulated with respect to the corresponding density functions, i.e., $f_{Q_{O\mid X}}(O,X)/f_{P_{O\mid X}}(O,X)=1$ almost everywhere in $P_{O\mid X}\times Q_{X}$.
Assumption 1 states that the support of $Q_{X}$ is within the support of $P_{X}$. Assumption 2 enables replacing $Q_{O\mid X}$ with $P_{O\mid X}$, which can be identified from the “full data” $(O,X)$ in the source population $P_{O,X}$. Under Assumptions 1 and 2, the target estimand can be identified as $\psi(Q_{X}\times P_{O\mid X})$ and some examples of target estimands are listed below.
@@
Suppose we are interested in the mean of $O$ in the target distribution, denoted as $\psi^{{\textnormal{mean}}}(Q_{O,X})=\textnormal{E}_{Q_{O}}(O)$. Under Assumptions 1 and 2, the mean is identified via
$\psi^{{\textnormal{mean}}}(Q_{O,X})=\textnormal{E}_{Q_{O}}(O)=$ $\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{Q_{O\mid X}}(O\mid X)\right]=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}(O\mid X)\right].$
@@
Suppose we are interested in the median of $O$ in the target distribution, denoted as $\psi^{{\textnormal{med}}}(Q_{O,X})=F_{Q_{O}}^{-1}(1/2)$, and $O$ is continuous. Under Assumptions 1 and 2, the median is identified as the solution to the following equation:
@@
Example 3 (Z-Estimand).
Suppose $\psi(Q_{O,X})$ is defined as the solution to
@@
where $s(O,X,\cdot)$ is a user-specified function, usually a score function of the same dimension as $\psi$. Under Assumptions 1 and 2, the target estimand is identified as the solution to
$0=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{Q_{O\mid X}}\left\{s(O,X,\psi(Q_{O\mid X}\times Q_{X}))\mid X\right\}\right]=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{O\mid X}}\left\{s(O,X,\psi(Q_{O,X}^{0}))\mid X\right\}\right].$
@@
Consider a randomized experiment in the source population to measure the average treatment effect (ATE). Let $Y(a)$ be the potential outcome if, contrary to fact, a study unit was assigned to treatment $a\in\mathcal{A}\subset\mathbb{R}$ where $\mathcal{A}$ is a set of all possible treatment (e.g., $\mathcal{A}=\{0,1\}$) and let $X$ be pre-treatment covariates. The goal is to learn about the ATE in a new, target population based on (a) the randomized experiment in the source population and (b) the distribution of $X$ in the target population. Under Assumptions 1 and 2, the mean and the median of the potential outcome $Y(a)$ in the target population are identified using Examples 1 and 2:
$\psi^{{\textnormal{mean}}}(Q_{Y(a),X})=\mathbb{E}_{Q_{Y(a)}}\{Y(a)\}=\textnormal{E}_{Q_{X}}\left[\textnormal{E}_{P_{Y(a)\mid X}}\{Y(a)\mid X\}\right],\text{ and }\quad{}\frac{1}{2}=\int\int_{-\infty}^{\psi^{{\textnormal{med}}}}dP_{Y(a)\mid X}dQ_{X}.$
We remark that both equations involve potential outcomes (i.e., $P_{Y(a)\mid X}$), which is identified under a randomized experiment in the source population; see Section D in the Supplement for details.
@@ 2.2 Model for Sensitivity Analysis of Conditional Exchangeability
Suppose we suspect that conditional exchangeability (i.e., Assumption 2) is implausible and we wish to assess how the conclusion of the study may change if conditional exchangeability is violated. A sensitivity analysis addresses this question by supposing that there is a $\gamma$ violation” of conditional exchangeability and quantifying the downstream consequences of this violation. This section presents a model-based sensitivity analysis (Rosenbaum and Rubin, 1983a, ; Robins et al.,, 2000; Franks et al.,, 2020) for quantifying violations of conditional exchangeability based on exponential tilting.
Formally, for each $\gamma\in\mathbb{R}$, let $Q_{O\mid X}^{\gamma}(\cdot\mid X)$ be absolutely continuous with respect to $P_{O\mid X}(\cdot\mid X)$ almost everywhere $Q_{X}$. Suppose the corresponding densities $f_{Q_{O\mid X}^{\gamma}}(O,X)$ and $f_{P_{O\mid X}}(O,X)$ satisfy the following relationship:
@@
$\dfrac{f_{Q_{O\mid X}^{\gamma}}(O,X)}{f_{P_{O\mid X}}(O,X)}\propto\exp(\gamma\cdot O)\text{, almost everywhere in }P_{O\mid X}\times Q_{X}.$ (3)
The notation $\propto$” means “proportional to” and the normalizing constant satisfies $\int\exp(\gamma O)dP_{O\mid X}(O\mid X)<\infty$ almost everywhere $Q_{X}$.
@@
The term $\gamma$ is often referred to as the sensitivity parameter and it measures the difference between $Q_{O\mid X}$ and $P_{O\mid X}$. If $\gamma=0$, the sensitivity model (3) reduces to Assumption 2 where the two distributions are identical, i.e., $Q_{O\mid X}^{0}=P_{O\mid X}$ almost everywhere in $Q_{X}$. As $\gamma$ moves away from zero, the difference between $Q_{O\mid X}$ and $P_{O\mid X}$ becomes larger and conditional exchangeability is violated by a larger amount.
@@
We make some remarks about the sensitivity model (3). First, model (3) was proposed by Scharfstein et al., (1999) and Robins et al., (2000) as a non-parametric (just) identified model for describing selection bias in missing data and has been used by several others (Rotnitzky et al.,, 2001; Birmingham et al.,, 2003; Troxel et al.,, 2004; Linero and Daniels,, 2018; Franks et al.,, 2020; Nabi et al.,, 2024; Dahabreh et al.,, 2022; Miao et al.,, 2024). Second, model (3) can be reformulated as a selection model and under some assumptions, the sensitivity parameter $\gamma$ can be reparameterized to pseudo-$R^{2}$ (Franks et al.,, 2020). Third, model (3) can be extended so that $\gamma$ depends on $X$ and $O$ or the exponential tilting term can be replaced with a non-negative tilting term (see Section 4.3). Fourth, an important caveat of model (3) is the non-collapsibility of the model with respect to $X$ as the model implies a logistic selection model; see Section 7 of Scharfstein et al., (1999). Fifth, model (3) differs from a “bound-based” sensitivity analysis (e.g., Rosenbaum, (1987); Tan, (2006)) and Remark 2 discusses an inherent difficulty in defining SLOPE with such models. In particular, model (3) (i) makes our proposed measure SLOPE tractable in terms of having a unique tangent curve (see Section 4.3), (ii) has an analytic connection to the influence function (see Section 4.1), (iii) posits no testable implications on the data (Franks et al.,, 2020), and most importantly, (iv) leads to simple and empirically validated principles about designing robust study designs for generalizations (see Sections 3.2, 3.3 and 6).
@@ 3.1 Definition and Basic Properties
Under Assumption 1 and the sensitivity model (3), with a chosen value of the sensitivity parameter $\gamma$, the target estimand can be identified via $\psi(Q_{O,X}^{\gamma})$ where $Q_{O,X}^{\gamma}=Q_{O\mid X}^{\gamma}\times Q_{X}$ represents the joint distribution of $Q_{O\mid X}^{\gamma}$ induced from (3) and $Q_{X}$. But, $\gamma$ is never known in practice because it represents the magnitude of violating conditional exchangeability (i.e., Assumption 2). Instead, sensitivity analysis seeks to understand how $\psi(Q_{O,X}^{\gamma})$ changes from $\gamma=0$ (i.e., when Assumption 2 holds) to $\gamma\neq 0$ (i.e., when Assumption 2 doesn’t hold). This is typically done by presenting a table or a plot of $\psi(Q_{O,X}^{\gamma})$ for a plausible range of $\gamma$ with the range determined by domain knowledge (Scharfstein et al.,, 1999; Rotnitzky et al.,, 2001; Nabi et al.,, 2024), benchmarking (Huang,, 2024) or calibration (Miao et al.,, 2024).
[…]
@@
The sensitivity to local perturbation from exchangeability (SLOPE) of a target functional/estimand $\psi$ with respect to the sensitivity model in (3) is defined as
@@
A large magnitude of SLOPE means that the target estimand $\psi$ will change more drastically if conditional exchangeability is slightly violated (i.e., $\gamma\to 0$). In contrast, a small magnitude of SLOPE means that the target estimand will change less drastically. Note that the magnitude of SLOPE is the absolute value of SLOPE when the target estimand $\psi$ is a scalar. When $\psi$ is a vector, the magnitude of SLOPE corresponds to the researcher’s choice of measuring the magnitude of vectors (e.g., $\ell_{2}$ norm). As remarked in Section 2.1, for expositional purposes, we focus on a scalar $\psi$, and thereby a scalar SLOPE, but our results hold for low-dimensional, vector-valued $\psi$.
When comparing the magnitudes of SLOPE, investigators should note that the unit of SLOPE inherits the unit of the target estimand, which is often defined with scientifically meaningful units. For instance, for the mean $\psi^{{\textnormal{mean}}}$ and the median $\psi^{{\textnormal{med}}}$ in Examples 1 and 2, respectively, the units of SLOPE for both estimands are the unit of $O$ and the two SLOPEs have identical units. If the investigator wishes to change the units of SLOPE, including a unit-less SLOPE, a simple approach is to change the units of the target estimand. More broadly, we recommend interpreting the magnitude of SLOPE with the same caution used for interpreting the magnitude of regression coefficients where the units of the regression coefficients inherit their underlying units from the data.
A key property of SLOPE is that it does not depend on any particular estimation procedure. Instead, SLOPE measures an intrinsic property about the robustness of a study design and is determined by two design quantities: (a) the target estimand $\psi$, and (b) the target distribution under conditional exchangeability ($Q_{O,X}^{0}=P_{O\mid X}\times Q_{X}$). For researchers, these choices roughly correspond to answering two questions: (a) “what quantity do I want to study?” and (b) “which dataset should I use to study the quantity?”. Changing the answers to either question can alter the value of SLOPE. Consequently, SLOPE can help researchers pick a robust study design by selecting an estimand (e.g., the mean or the median of $O$ in the target distribution), target distribution (e.g., $Q_{X}$), or the source distribution (e.g., $P_{O,X}$) that leads to a lower magnitude of SLOPE; see Sections 3.2 and 5 for illustrations.
When communicating the meaning of SLOPE, some researchers may find it useful to interpret SLOPE as a measure of the “first-order change” of the estimand when conditional exchangeability is violated. Specifically, a first-order Taylor expansion of $\psi$ yields
@@
The Taylor expansion suggests that for a small $\gamma$ that is near zero, SLOPE provides an accurate measure of the change in $\psi$ when conditional exchangeability is violated. When $\gamma$ is large in magnitude, SLOPE may still provide some intuition about the change in $\psi$, with the usual limitations of first-order linear approximations. We remark, however, that SLOPE cannot identify the bias of an estimator of $\psi$ from violating conditional exchangeability since, as mentioned in Section 2.2, $\gamma$ and $Q_{O,X}^{\gamma}$ are not identifiable.
@@
Finally, we remark that SLOPE does not always exist for every target estimand. For example, if $\psi$ is the sign of the mean of $O$ in the target population, the limit in Definition 1 may not exist when the sign changes near $\gamma=0$. One general condition for SLOPE to exist is to satisfy the conditions for the chain rule under Hadamard differentiability; see Section A of the Supplement for details. For Z-estimands in Example 3, a sufficient condition for the existence of SLOPE is to impose smoothness and boundedness conditions on $s$; note that these conditions are common to establish consistency of Z-estimators.
Condition 1 (Existence of SLOPE for Z-Estimands).
@@
(i) $\textnormal{E}_{Q_{O\mid X}^{\gamma}}[s(O,X,\psi(Q_{O,X}^{0}))]$ is bounded for $\gamma$ in a neighborhood of zero, and $\textnormal{E}_{Q_{O,X}^{0}}\left[s(O,X,\psi(Q_{O,X}^{0}))\left\{O-\mu(X)\right\}\right]$ exists where $\mu(X)=\textnormal{E}_{P_{O\mid X}}[O\mid X]$; (ii) $s(O,X,\cdot)$ is differentiable almost everywhere with the derivative $\dot{s}(O,X,\cdot)$; and (iii) $\textnormal{E}_{Q_{O,X}^{0}}\{\dot{s}(O,X,\psi(Q_{O,X}^{0}))\}$ exists and is non-singular.
3.2 Example: SLOPE for the Mean and Robust Study Designs
This section has two main goals. The first is to show that how researchers can derive SLOPE for a given $\psi(Q_{O,X}^{\gamma})$ using basic calculus. The second is to illustrate how SLOPE can yield useful insights about robust study designs for generalization.
To begin, consider the mean of $O$ in the target population, i.e., $\psi^{{\textnormal{mean}}}$ in Example 1 where $\psi^{\textnormal{mean}}(Q_{O,X})=\textnormal{E}_{Q_{O}}(O)$. From Section 3.1, for a given $\gamma$, the sensitivity model (3) implies the following equality:
$\psi^{{\textnormal{mean}}}(Q_{O,X}^{\gamma})=\textnormal{E}_{Q_{O,X}^{\gamma}}(O)=\textnormal{E}_{Q_{X}}\left\{\textnormal{E}_{Q_{O\mid X}^{\gamma}}\left(O\mid X\right)\right\}=\textnormal{E}_{Q_{X}}\left[\dfrac{\textnormal{E}_{P_{O\mid X}}\left\{O\exp(\gamma O)\mid X\right\}}{\textnormal{E}_{P_{O\mid X}}\left\{\exp(\gamma O)\mid X\right\}}\right].$